Biochemical Corp. requires $720,000 in financing over the next three years. The frm can borrow the funds for three years at 10.20 percent interest per year. The CEO decides to do a forecast and predicts that if she utilizes short term financing instead, she will pay 8.50 percent interest in the first year, 12.90 percent interest in the second year, and 9.75 percent interest in the third year. Assume interest is pald in full at the end of each year. a. Determine the total interest cost under each plan. Interest Cost Long-term fixed-rale Shorn term variable-rate b. Which plan is less costly? Short-tem variable-rate plan O Long-term fixed-rate plarn
Answer #1

Please provide rating...

Loan amount = 720000
Answer a
Long term Fixed rate interest
For 3 year =720000*10.2%*3 220320
Sort term variable rate interest
Year -1 =720000*8.5% 61200
Year -2 =720000*12.9% 92880
Year -3 =720000*9.75% 70200
Total interest cost 224280
Long term Fixed rate interest 220320
Sort term variable rate interest 224280
Answer b.
Long term Fixed rate plan cost less